OpenStudy (anonymous):

To which of the given expressions is ln 1/5 equivalent? A. ln 5 − ln 1 B.1/5 C.1/ln 5 D. ln 1 − ln 5 E. ln 5 F. − ln 5 G. ln 5 + ln 1 H. None of these

OpenStudy (anonymous):

\[\ln\frac{ a }{ b} = \ln a - \ln b\]

OpenStudy (anonymous):

ln1/5 = ln 1 - ln 5 since, ln 1 = 0 ln1/5 = -ln 5